OpenStudy (xmoses1):

Which one of the following equations illustrates the commutative property? 8 + 3 = 4 + 7 If 3 + 5 = a, then a = 3 + 5 2 + b = b + 2 3(2 + 5) = 6 + 15

OpenStudy (anonymous):

I THINK ITS 2 + b = b + 2

OpenStudy (neer2890):

. A binary operation * on a set S is called commutative if: x * y = y * x {for all }x,y in S

OpenStudy (neer2890):

so it's 2+b=b+2

OpenStudy (xmoses1):

Hey thanks guys :) Could you medal each other? I want to medal both of you :p